### v1

# Continuous Monitoring via Repeated Significance

Abstract
Requiring statistical significance at multiple interim analyses to declare a statistically significant result for an AB test allows less stringent requirements for significance at each interim analysis. Repeated repeated significance competes well with methods built on assumptions about the test – assumptions that may be impossible to evaluate a priori and may require extra data to evaluate empirically.
Instead, requiring repeated significance allows the data itself to prove directly that the required results are not due to chance alone. We explain how to apply tests with repeated significance to continuously monitor unbounded tests – tests that do not have an a priori bound on running time or number of observations. We show that it is impossible to maintain a constant requirement for significance for unbounded tests, but that we can come arbitrarily close to that goal.
I Introduction
Recently, there has been renewed interest in group sequential methods [1, 2, 3, 4, 5, 6], or interim analysis methods [7, 8, 9] for long-running AB tests, which may lack a priori limits on run time or number of observations. We refer to such tests as unbounded AB tests. Methods that control type 1 error for such tests include always-valid bounds and continuous monitoring methods [10, 11, 12, 13, 14].
Most group sequential methods use some form of $\alpha$ spending – partitioning a type 1 error budget over the interim analyses [15, 6]. Adapting them to unbounded tests requires partitioning the error budget over an unbounded number of interim analyses. Dividing the budget more finely (while controlling type 1 error) makes conditions for ending the test more stringent, increasing test time and expense and increasing the likelihood of type 2 error.
To counter those effects, always-valid bounds and continuous monitoring methods often rely on correlation between $p$-values over sequences of interim analyses. Those correlations follow from a combination of properties of specific statistics and requirements for the observations. For example, a running average over observations combined with a limit on the value of each observation implies that the running average will have limited change as the number of observations grows large.
Another method to avoid extremely stringent conditions for large numbers of interim analyses is to require repeated statistically significant results [16] over multiple interim analyses to declare a statistically significant result for the AB test as a whole. Requiring repetition makes each interim significance requirement less stringent by multiplying the $p$-value required for significance by the number of repeats required – a consequence of nearly uniform validation [17, 18].
Compared to other methods, requiring repetition removes the need to know a priori, prove, or empirically evaluate whether side conditions on the data hold, because it does not rely on correlations among $p$-values over sequences of interim analyses. Instead, it allows (and requires) the data to prove itself. This makes it simple and widely applicable, yet effective, as has been shown for bounded tests [16], and as we show in this paper for unbounded tests.
Section II reviews the concepts of $\alpha$ spending for sequential testing, nearly uniform validation, and requiring repeated significance. Section III shows how to apply those concepts to unbounded tests. Section IV explores using $\alpha$ spending based on geometric series to produce test plans that control type 1 error. Section V shows that, for unbounded tests, the $p$-values required for statistically significant results must grow eventually, but we can delay that using $\alpha$ spending based on $p$-series with exponents near one. Section VI concludes with a discussion of possibilities for future work.
II Preliminaries
For a statistical test, a null hypothesis is an assertion about the (unknown) out-of-sample statistics or underlying distribution that generates the observations in the test. For a pre-specified metric measured on the in-sample observations, a $p$-value is the probability that the metric takes a value at least as extreme as the measured value under the null hypothesis.
A type 1 error occurs when a statistical test rejects the null hypothesis but the null hypothesis holds – a false positive. In test planning, we specify an acceptable upper bound for the probability of type 1 error. Call that bound $\alpha$. If the test rejects the null hypothesis, then we say it is rejected at confidence level $1-\alpha$.
A simple test plan specifies a metric and $\alpha$. After all in-sample observations are collected, the test measures the metric and produces a $p$-value based on the measured value. If $p\leq\alpha$, then the null hypothesis is rejected.
For example, in an AB test, the null hypothesis may be that both treatments A and B (one of which may be the control) have the same average for some metric. If the $p$-value for the test is below a specified $\alpha$, then we reject the null hypothesis, asserting that the difference in treatments has a statistically significant impact on the metric, with confidence $1-\alpha$.
For many AB tests, it is useful to allow early stopping – ending the test with a statistically significant result before collecting all planned observations. In tests of user interfaces or advertisements on apps and web pages, this can reduce time to market and save money. In medical tests, it can save lives by delivering useful treatments to the control group as well as the treatment group as soon as there is a positive result.
Early stopping can create the risk of data dredging or $p$-hacking [19, 20, 21] in time, by stopping when a metric is positive but continuing when it is negative. That can introduce large gaps between in-sample and out-of-sample results, due to a selection effect [22, 23, 24, 7].
In general, stopping when the $p$-value for the data collected from the start of the test to an interim analysis is less than $\alpha$ fails to bound the type 1 error probability by $\alpha$. If the probability of type 1 error at each of $d$ interim analyses is at most $\alpha$, then the probability of type 1 error at any interim analysis may be as high as $d\alpha$. (In the worst case, each interim type 1 error probability is $\alpha$, and those events are disjoint.)
A technique to avoid this problem is called $\alpha$ spending – partitioning $\alpha$ into $\alpha_{1},\ldots,\alpha_{d}$ that sum to $\alpha$, and allowing early stopping when the $p$-value for interim analysis $k$ (call it $p_{k}$) has $p_{k}\leq\alpha_{k}$. In this case, the probability of type 1 error at each interim analysis $k$ is at most $\alpha_{k}$, so the overall probability of a type 1 error is at most $\alpha$, by the sum bound on the union of probabilities, also known as Boole’s inequality [25, 26]. (Dividing evenly: $\alpha_{k}=\frac{\alpha}{d}$ is known as a Bonferroni correction [27].) The method of $\alpha$ spending has a long history in group sequential testing [1, 28, 2, 3, 4, 5, 6], especially for medical trials.
As $\alpha$ is partitioned more finely over more interim analyses, the $p$-value requirements for statistical significance become more stringent. Requiring repeated significance in order to stop the test is a way to counteract this effect, because it allows less stringent requirements for statistical significance while maintaining the $\alpha$ bound on probability of type 1 error. If we require $p_{k}\leq\delta_{k}$ for at least $r$ interim analyses before ruling out the null hypothesis and ending the test, then the probability of type 1 error is at most $\frac{\delta_{1}+\ldots+\delta_{d}}{r}$, because the probability of type 1 error is then maximized by having $p_{k}\leq\delta_{k}$ co-occur for exactly $r$ different $k$ values in $1,\ldots,d$ when it occurs for any $k$.
Requiring $r$ failures for a type 1 error allows us to divide the sum bound by $r$ (see Figure 1), or, equivalently to multiply each budgeted $\alpha_{k}$ by $r$ to get $p$-value requirements for interim analyses, since $\alpha_{1}+\ldots+\alpha_{d}=\alpha$ implies that if we set $p$-value requirements $\delta_{k}=\alpha_{k}r$, then $\frac{\delta_{1}+\ldots+\delta_{d}}{r}=\alpha$.
Figure 1: Requiring Repetition. For $r=3$ required repeats over $d=6$ interim analyses each with interim analysis type 1 error probability $\delta$, the worst-case probability of type 1 error for the entire test is $\frac{6\delta}{3}=2\delta$, since 3 interim analyses must all have type 1 errors for the entire test to have it. (If six carpets each cover $\delta$ area, then at most $2\delta$ area can be covered at least three deep.) (Figure from [18].)
An earlier paper [16] applies this insight to develop repetition-based test plans for tests with a priori limits on the number $d$ of interim analyses. For that case, it is possible, for example, to spend $\alpha$ evenly over interim analyses: $\alpha_{k}=\frac{\alpha}{d}$ for all $k$ in $1,\ldots,d$. For unbounded $d$, that is impossible. This paper focuses on methods to spend $\alpha$ and select the number of required repeats for unbounded tests.
III Unbounded Tests
Following the notation from [16], we will refer to interim analyses as decision points. For each decision point $t$, let $\delta_{t}$ be the upper bound for the $p$-value at decision point $t$ to be considered significant, and let $r_{t}$ be the number of significant $p$-values over decision points $1$ to $t$ required to end the test and declare a significant result.
As an example, suppose that for each decision point $t$, if at least half of the decision points $k$ from $1$ to $t$ have $p_{k}\leq\delta_{k}$, where $p_{k}$ is the $p$-value at decision point $k$, then we end the test and declare a significant result – rejecting the null hypothesis. Then $r_{t}=\lceil\frac{t}{2}\rceil$ for all $t\geq 1$.
Theorem 1.
Algorithm 1 returns the greatest possible type 1 error rate $\alpha$ given $\delta_{1},\delta_{2},\ldots$ and $r_{1}\leq r_{2}\leq\ldots$, for a test that stops at the first decision point $t$ at which at least $r_{t}$ of the decision points $k$ from $1$ to $t$ each have $p_{k}\leq\delta_{k}$, with $p_{k}$ the $p$-value at decision point $k$.
Proof.
Each decision point $k$ has probability at most $\delta_{k}$ of false significance – having its $p$-value at most $\delta_{k}$ even though the null hypothesis holds. The event that a type 1 error results from stopping at decision point $t$ requires at least $r_{t}$ cases of false significance over $k$ from $1$ to $t$. The type 1 error probability for the test, $\alpha$, is the probability of the union of those events over $t$ in $1,2,\ldots$. So a joint distribution of $\delta_{k}$ false significance probabilities that maximizes the type 1 error probability for the test should:
  1. 1.
    Use false significances as efficiently as possible by having as few of them occur simultaneously as needed to create type 1 error events, and
  2. 2.
    Have those events be disjoint, to maximize the probability of their union.
Algorithm 1 finds $\alpha$ for such a distribution. It uses false significances efficiently by collecting them into simultaneous sets of $r_{t}$ false significances for the earliest $t$ possible – that greedy approach uses as few false significances as possible to create type 1 error events because $r_{1}\leq r_{2}\leq\ldots$, so at least as many false significances would be required to form a type 1 error event later. The algorithm sums the probability of those type 1 error events into $\alpha$, so it calculates the probability of the union of those events for a distribution in which the events are disjoint. ∎
1 Input: $\delta_{1},\delta_{2},\ldots$ and $r_{1},r_{2},\ldots$
2 Initialize: $L\leftarrow[\hbox{ }]$, $\alpha\leftarrow 0$
3 for $t\in 1,2\ldots$ do
4       Append $\delta_{t}$ to $L$
5      if $L$ has $r_{t}$ entries then
6             $m\leftarrow\min L$
7             $\alpha\leftarrow\alpha+m$
8             Subtract $m$ from each entry in $L$
9             Remove all zero-valued entries from $L$
return $\min(\alpha,1)$
Algorithm 1 Bound Type 1 Error Probability
Algorithm 1 applies to the case of requiring a fixed fraction of decision points up to each decision point $t$ to have significant results in order to stop the test, because then $r_{1}\leq r_{2}\leq\ldots$. It also applies to unbounded tests: tests that do not have a priori limits on the number of decision points.
As an example, again consider requiring at least half of the decision points up to each decision point $t$ to have significant results in order to stop at decision point $t$. Then $r_{t}=\lceil\frac{t}{2}\rceil$. Setting
$\delta_{t}=\frac{\alpha}{2^{\lceil\lg t\rceil+1}}$ (1)
($\lg$ is the base 2 logarithm), the algorithm gathers $\delta_{t}$ values to have type 1 errors at $t$ values that are powers of two:
$\boxed{\frac{\alpha}{2}},\boxed{\frac{\alpha}{4}},\boxed{\frac{\alpha}{8},\frac{\alpha}{8}},\boxed{\frac{\alpha}{16},\frac{\alpha}{16},\frac{\alpha}{16},\frac{\alpha}{16}},\ldots$ (2)
Since $r_{1}=r_{2}=1$, the first two $\delta_{t}$ values add directly to $\alpha$. Since $r_{3}=r_{4}=2$, the algorithm gathers $\delta_{3}$ and $\delta_{4}$ to make one type 1 error event. Then it gathers the next four $\delta_{t}$ values, because $r_{8}=4$. Thereafter, it gathers the next $8,16,\ldots$ values.
The type one error probability is at most $\alpha$, since
$\frac{\alpha}{2}+\frac{\alpha}{4}+\frac{\alpha}{8}+\frac{\alpha}{16}+\dots=\alpha$ (3)
Stopping the test at decision point $t$ requires
$p_{k}\leq\frac{\alpha}{2^{\lceil\lg k\rceil+1}}$ (4)
for at least half of the decision points $k$ in 1 to $t$.
A non-algorithmic result follows from Theorem 1:
Corollary 1.
Given $\delta_{1},\delta_{2},\ldots$ and $r_{1}\leq r_{2}\leq\ldots$, for a test that stops at the first decision point $t$ at which at least $r_{t}$ of the decision points $k$ from $1$ to $t$ each have $p_{k}\leq\delta_{k}$, with $p_{k}$ the $p$-value at decision point $k$, an upper bound on the type 1 error probability is
$\sum_{t}\frac{\delta_{t}}{r_{t}}$ (5)
Proof.
In Algorithm 1, appending $\delta_{t}$ to $L$ adds $\delta_{t}$ to the sum of the entries in $L$. Later, for zero or more values $k\geq t$, the algorithm’s if-then block adds $\frac{1}{r_{k}}$ as much to $\alpha$ as it subtracts from the sum of entries in $L$$m$ added to $\alpha$ versus $r_{k}m$ subtracted from the sum over $L$. Since $r_{1}\leq r_{2}\leq\ldots$, $\frac{\delta_{t}}{r_{k}}\leq\frac{\delta_{t}}{r_{t}}$. ∎
We can restate Corollary 1 from an $\alpha$-spending point of view:
Corollary 2.
Given $r_{1}\leq r_{2}\leq\ldots$ and a set of nonnegative values $\alpha_{t}$ such that
$\sum_{t}\alpha_{t}=\alpha$ (6)
if a test stops at the first decision point $t$ at which at least $r_{t}$ of the decision points up to $t$ have $p$-values with:
$p\leq\alpha_{t}r_{t},$ (7)
then the test has type 1 error probability at most $\alpha$.
Proof.
Let $\delta_{t}=\alpha_{t}r_{t}$ and apply Corollary 1. ∎
For bounded tests – ones that have an priori limit $d$ on the number of decision points – it is possible to use uniform $\alpha$ spending, with all $\alpha_{k}=\frac{\alpha}{d}$, where $d$ is the limit on the number of decision points. Similarly, it is possible to budget $\alpha$ so that the $p$-value requirement for significance, $\delta_{t}=\alpha_{t}r_{t}$, is equal for all $t\in\left\{1,\ldots,d\right\}$.
For unbounded tests, geometric spending allocates the same fraction of the remaining $\alpha$ budget to each successive decision point: select a “withdrawal rate” $10<w<$ 1 and set $\alpha_{1}=w\alpha$, $\alpha_{2}=w(1-w)\alpha$, and in general $\alpha_{t}=w(1-w)^{t-1}\alpha$. Decreasing $w$ decreases early $\alpha_{t}$ values, saving more of $\alpha$ for later ones. Geometric spending works for unbounded tests, since the sum of $\alpha_{t}$ values are the sum of a geometric series that sums to $\alpha$ in the limit.
As an example applying Corollary 2 to geometric $\alpha$ spending, if the test plan is to require at least $r_{t}=\lceil ut\rceil$ significant results over decision points 1 to $t$ to stop at decision point $t$, then significance for each decision point $t$ requires
$p_{t}\leq w(1-w)^{t-1}\alpha\lceil ut\rceil$ (8)
for a value $10<w<$ 1 specified as part of the test plan.
IV Geometric $\alpha$-Spending
Figures 2 to 4 show $Z$ scores required for significance at each decision point for an unbounded test that allows stopping at the first decision point $t$ with a fraction at least $u$ of the decision points 1 to $t$ having $p$ values low enough for significance at their decision points, with probability of type 1 error at most $\alpha=0.05$. We plot $Z$ scores rather than $p$ values for ease of presentation, since $Z$ scores are familiar to many practitioners and have a 1-1 correspondence with $p$ values. The plotted $Z$ scores are the number of standard deviations from the mean of a standard normal such that the left and right tails together have probability $p$. (For comparison, a test with a single decision point that spends all of $\alpha=0.05$ on that decision point would require $Z\approx 1.96$ for $p=0.05$.)
The figures show that decreasing the “withdrawal” rate increases the number of decision points $t$ before reaching the minimum $Z$ score required for significance and decreases the rate of increase after reaching the minimum. Decreasing $w$ maintains more of the $\alpha$ budget for later decision points.
Together, the figures show that requiring more repetition makes requirements for significance at each decision point less stringent: as $u$ increases from $0.05$ to $0.10$ then to $0.5$ in Figures 2 to 4, the curves drop to lower $Z$ scores. (The exception is the first few values of $t$ on the left, since they require the same number of repetitions: $\lceil ut\rceil=1$ until $t>\frac{1}{u}$.)
Figure 2: For $u=0.05$, $\mathbf{Z}$ score for significance at each decision point. Requiring 5% of decision points to have significant results before ending the test makes the minimum required $Z$ scores about 3.3. A smaller “withdrawal” rate $w$ increases the number of decision points $t$ before reaching the minimum.
Figure 3: For $u=0.1$, $\mathbf{Z}$ score for significance at each decision point. Requiring 10% of decision points to have significant results before ending the test makes the minimum required $Z$ scores about 3.1 – lowering the curves from Figure LABEL:fiz_zgeo005 for $u=0.05$ by requiring more repetition.
Figure 4: For $u=0.5$, $\mathbf{Z}$ score for significance at each decision point. Requiring 50% of decision points to have significant results before ending the test makes the minimum required $Z$ scores about 2.6, compared to about 3.1 for $u=0.1$ and 3.3 for $u=0.05$ (as shown in Figures 3 and 2).
Figures 5 and 6 show results for larger numbers of decision points, both with $u=0.1$, as in Figure 3 (all with $\alpha=0.05$). To be clear, all those figures show results for unbounded tests, so they all have unlimited numbers of decision points, what is different is the number of initial decision points shown in each plot: ten thousand, a million, then a billion. Note that, after adjusting the values of $w$ in inverse proportion to the number of decision points plotted, the resulting curves are very similar, including their minimum $Z$ scores. So, by selecting $w$ appropriately, we can hold $u$ constant and have nearly the same curve, but stretched horizontally, with the same test effectiveness in terms of the $p$-values required for significant results. That said, increasing $t$ while holding $u$ constant maintains the same repetition requirement in terms of fraction of decision points, but increases it in absolute terms.
Figure 5: For $u=0.1$ and over the first million decision points, $\mathbf{Z}$ scores for significance. The “withdrawal” rates $w$ in this figure are one hundredth those in Figure 3, which is over the first ten thousand decision points – one hundredth of a million. The result is similar curves, with similar minimum $Z$ scores required for significance.
Figure 6: For $u=0.1$ and over the first billion decision points, $\mathbf{Z}$ scores for significance. The “withdrawal” rates $w$ in this figure are one thousandth those in Figure 5, which is over the million decision points, giving similar $Z$ score curves.
Figure 7 compares using geometric $\alpha$ spending with repetition requirement $u=0.2$ to an always-valid method based on autocorrelation among averages [14] (their Equation 8, page 8). The $Z$ score required for significance for that method is
$\sqrt{\frac{2(t\rho^{2}+1)}{t\rho^{2}}\ln\left(\frac{\sqrt{t\rho^{2}+1}}{\alpha}\right)},$ (9)
with $\rho$ a parameter that controls the $t$ value at which the $Z$ score reaches a minimum, like $w$ for our method. Both methods apply to unbounded tests.
For $u=0.2$, repetition generally produces lower $Z$ score requirements. However, requiring repetition with $u=0.2$ may require a test to run for up to 20% longer, if it remains significant once it becomes significant for some $t$ value. Of course, due to the less stringent $Z$ score requirement, requiring repetition may also result in earlier stopping than for the other always-valid method. For $u\approx 0.15$, both methods have similar $Z$ score minimums over $t$.
Figure 7: Comparison to another always-valid method. For $u=0.2$, requiring repetition gives less stringent requirements for significance than a method based on autocorrelation for averages [14].
V Nearly Impossible is Possible
Figures 2 to 6 all have $Z$ score curves that increase after reaching a minimum, prompting the question: “Is there a way to spend $\alpha$ that gives a flat curve for an unbounded test, keeping the required $p$-value from shrinking to zero, or, equivalently, the required $Z$ score from becoming unbounded as $t\rightarrow\infty$?”
To answer that question, first consider that, in general, for an unbounded test we require
$\sum_{t=1}^{\infty}\alpha_{t}\leq\alpha$ (10)
to avoid over-spending $\alpha$. If we let
$\alpha_{t}=x_{t}\alpha,$ (11)
then we require
$\sum_{t=1}^{\infty}x_{t}\leq 1,$ (12)
so we need $x_{t}$ to be a convergent series.
Now we show that the answer to the question is a negative of the form: “It is not quite possible, but nearly so.”
Theorem 2.
It is impossible to have all $p$-value requirements $\delta_{t}$ given by Corollary 2 equal for an unbounded test if $r_{t}=\lceil ut\rceil$ for a constant $u>0$.
Proof.
Recall that
$\delta_{t}=\alpha_{t}r_{t}=x_{t}\alpha\lceil ut\rceil.$ (13)
Suppose these values are all equal to a constant $c$. Then, for all $t\in\left\{1,2,\ldots\right\}$,
$x_{t}=\frac{c}{\alpha\lceil ut\rceil}>\frac{c}{\alpha(ut+1)}$ (14)
$=\frac{c}{\alpha u}\left(\frac{1}{t+\frac{1}{u}}\right).$ (15)
We require the sum $x_{1}+x_{2}+\dots\leq 1$, but the sum of these terms diverges. Informally, it diverges because the terms are on the order of harmonic series terms: $\frac{1}{t}$, and the harmonic series diverges.
More formally, note that:
$\frac{1}{t+\frac{1}{u}}=\frac{1}{t}-\frac{\frac{1}{u}}{t(t+\frac{1}{u})}.$ (16)
So
$\sum_{t=1}^{\infty}x_{t}>$ (17)
$\frac{c}{\alpha u}\left[\sum_{t=1}^{\infty}\frac{1}{t}-\frac{1}{u}\sum_{t=1}^{\infty}\frac{1}{t(t+\frac{1}{u})}\right]$ (18)
The second sum is less than
$\sum_{t=1}^{\infty}\frac{1}{t^{2}}.$ (19)
We know that the $p$-series
$\sum_{t=1}^{\infty}\frac{1}{t^{v}}$ (20)
diverges for $0\leq v\leq 1$ and converges for $v>1$. So in Expression 18, the first sum diverges and the second converges, implying that the expression diverges. ∎
In short, we cannot have $x_{t}$ proportional to $\frac{1}{t}$ because its $p$-series diverges. However, we can use $x_{t}$ proportional to any $\frac{1}{t^{v}}$ for any $v>1$, since its $p$-series converges. So, in a sense, Theorem 2 is the most nearly possible impossible result possible.
Since
$\sum_{t=1}^{\infty}\frac{1}{t^{v}}=\zeta(v),$ (21)
where $\zeta()$ is the Riemann zeta function, we can set
$\forall t:\hbox{ }x_{t}=\frac{1}{\zeta(v)t^{v}}$ (22)
for any $v>1$. As $v\rightarrow 1$, we flatten the $\delta_{t}$ curve, but $\zeta(v)$ increases, making the initial $p$-value requirement more stringent. For example, $\zeta(2)=\frac{\pi^{2}}{6}$, so we can set
$x_{t}=\frac{6}{\pi^{2}t^{2}},$ (23)
and get
$\delta_{t}=\frac{6}{\pi^{2}t^{2}}\alpha\lceil ut\rceil\geq\frac{6\alpha u}{\pi^{2}t}.$ (24)
Figures 8 shows the $Z$ scores required for significance, using $p$-series $\alpha$ spending with different exponents $v$, with
$x_{t}=\frac{1}{\zeta(v)t^{v}},$ (25)
so
$\delta_{t}=\frac{\alpha\lceil ut\rceil}{\zeta(v)t^{v}}.$ (26)
As $v\rightarrow 1$, the curve of required $Z$ scores becomes flatter. Figure 9 shows the same results in detail for the first hundred decision points. Note that as $v\rightarrow 1$, the $Z$ scores decrease for the initial few decision points, and the curves are not very flat initially.
Figure 8: For $p$-series $\alpha$ spending, $\mathbf{Z}$ scores for significance. Using $p$-series spending, the curves flatten as $v$ approaches $1$ from the left, though we cannot set $v$ to exactly one without overspending $\alpha$, by Theorem 2. ($u=0.1$ and $\alpha=0.05$.)
Figure 9: Detail: for $p$-series $\alpha$ spending, $\mathbf{Z}$ scores for significance. This is the data from Figure 8, for the first hundred decision points. The curves are not flat for the first few decision points, so $\alpha$ is, in a sense, “overspent” initially.
To smooth the curve of $\delta_{t}$ over $t$, it is possible to remove the head of the $p$-series and use only its tail for $x_{t}$ values. Remove the first $s$ terms and let
$h(s,v)=\sum_{t=1}^{s}\frac{1}{t^{v}}$ (27)
be the head of the $p$-series consisting of the first $s$ terms. Removing that head yields
$x_{t}=\frac{1}{\left(\zeta(v)-h(s,v)\right)\left(t+s\right)^{v}},$ (28)
and so
$\delta_{t}=\frac{1}{\left(\zeta(v)-h(s,v)\right)\left(t+s\right)^{v}}\alpha\lceil ut\rceil.$ (29)
Figures 10 shows the results of removing some initial terms from the $p$-series to produce $x_{t}$. The curve for $s=0$ has no terms removed; it is the curve for $v=1.1$ from Figure 8. Note that removing more terms lowers the required $Z$ scores for significance for most of the $t$ values. Figure 11 shows the same data, but just for the first hundred decision points, where the ordering is reversed: removing more terms increases required $Z$ scores for the initial decision points.
Figure 10: Headless $p$-series $\alpha$ spending: $\mathbf{Z}$ scores for significance. For $p$-series $\alpha$ spending with $v=1.1$, $\alpha=0.05$, and $u=0.1$, removing more initial values from the $p$-series before starting to allocate the series values to $x_{t}$ lowers required $Z$ scores on the right.
Figure 11: Detail: headless $p$-series $\alpha$ spending: $\mathbf{Z}$ scores for significance. Plotting the data of Figure 10 shows that the lower $Z$ scores on the right of that plot are at the expense of raised $Z$ score requirements for the first few decision points.
VI Discussion
We have explored different ways to spend $\alpha$ with repetition requirements for unbounded tests, showing that it is possible to adjust the levels of significance required at each decision point by requiring more repetition. We have also shown that it is impossible to have a significance requirement that does not eventually become more stringent for an unbounded test, but we can stave that off for an arbitrary number of decision points by setting parameters appropriately.
An earlier paper [16] that focused on bounded tests discussed how to balance control of type 1 and type 2 error, by spending a portion of $\alpha$ on a single loose requirement for significance at the end of the test, which is reminiscent of traditional sequential testing strategies, such as the Haybittle-Peto boundary [28, 2]. It is a challenge to imagine something similar for unbounded tests, since their ending decision points are not known a priori. Perhaps spending a portion of $\alpha$ for loose requirements at a (possibly unbounded) subsequence of the decision points, either with or without a repetition requirement, would prove useful.
It would also be interesting to explore ways to combine repetition requirements with other approaches to continuous monitoring and always-valid bounds [10, 11, 12, 13, 14]. In some cases, it may be possible to improve both types of bounds by combining their strategies. The challenge is that non-repetition methods sometimes make use of side conditions that are proven empirically in some sense with repetition requirements – we need to analyze how these approaches overlap to understand how they can complement each other beyond mere duplication, if possible.